# Improving the Euclidean Diffusion Generation of Manifold Data by Mitigating Score Function Singularity

### v1

@@ 4.1 Niso-DM: Perturb data with non-isotropic noise
We propose a strategy to mitigate scale discrepancies by replacing the isotropic noise in the forward process of diffusion models with non-isotropic noise. The perturbed data is generated as $\tilde{x}_{t}=x+\sigma_{t}\epsilon_{1}+\sigma_{t}^{\alpha_{t}}N(x)\epsilon_{2}$, where $x\sim p_{0}(x)$, $\epsilon_{1}\sim\mathcal{N}(0,I_{n})$, $\epsilon_{2}\sim\mathcal{N}(0,I_{n-d})$, $\alpha_{t}\in(0,1)$, and $N(x)\in\mathbb{R}^{n\times(n-d)}$ is defined in Section 2.2. The term $\sigma_{t}^{\alpha_{t}}N(x)\epsilon_{2}$ represents an additional noise along the normal direction. The conditional distribution of the perturbed data is given by $p_{\sigma_{t}}(\tilde{x}|x)=\mathcal{N}(x,\Sigma_{\sigma_{t}}(x))$, where $\Sigma_{\sigma_{t}}(x)=\sigma_{t}^{2}I+\sigma_{t}^{2\alpha_{t}}N(x)N(x)^{T}$. The following Theorem establishes the relationship between $\nabla_{\tilde{x}}\log p_{\sigma_{t}}(\tilde{x})$ and $\nabla_{\tilde{x}}\log p_{0}(\tilde{x})$, as $\sigma_{t}\to 0$.
@@
Let $p_{\sigma}(\tilde{x})$ denote the distribution under non-isotropic perturbation, defined as:
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}e^{-V_{0}(x)}(2\pi)^{-\frac{n}{2}}(\det\Sigma_{\sigma}(x))^{-\frac{1}{2}}e^{-\frac{1}{2}(\tilde{x}-x)^{T}\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)}\mathrm{d}\sigma(x),$ (16)
where $\Sigma_{\sigma}(x)=\sigma^{2}I+\sigma^{2\alpha}N(x)N(x)^{T}$ and $\alpha\in(0,1)$. By making the same assumptions as in Theorem 3.1, and $M_{2}=\sup_{x\in\mathcal{M}}\max_{1\leq k,l,j,j^{\prime}\leq n}\left|\frac{\partial^{2}P_{kl}}{\partial x_{j}\partial x_{j^{\prime}}}(x)\right|<+\infty$, then we have the following results for $p_{\sigma}(\tilde{x})$:
@@
    $\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{x^{*}-\tilde{x}}{\sigma^{2\alpha}}\cdot\frac{1}{1+\sigma^{2-2\alpha}}+O(\sigma^{(1-2\alpha)\wedge 0}).$ (17)
@@
Eq. (17) shows that by introducing additional noise along the normal direction, the scale of the normal component is reduced from $O(1/\sigma^{2})$ to $O(1/\sigma^{2\alpha})$ pointwise. Eq. (18) states that this approach does not affect the characterization of the distribution on the manifold (i.e., the tangential component). As the noise level parameter $\sigma_{t}$ approaches zero, the score function of the perturbed distribution $p_{\sigma_{t}}$ on the manifold converges to the Riemannian score function of $p_{0}$. Therefore, by employing non-isotropic noise, we can still generate samples that adhere to the original distribution of the dataset.
@@ 4.2 Tango-DM: Learn only the tangential component of the score function
Recall that the projection matrix $P(x)$ can be continuously extended to the ambient space $\mathbb{R}^{n}$. Accordingly, we define $s^{\parallel}_{\theta}(x,t):=P(x)s_{\theta}(x,t)$ for $x\in\mathbb{R}^{n}$, which represents the tangential component of the parametrized vector field.
As discussed in (13), the loss function (4) can decomposed into two parts, $\ell_{\text{quad}}^{\parallel}$ and $\ell_{\text{quad}}^{\perp}$, and the singularity issue primarily arises from the normal component $\ell_{\text{quad}}^{\perp}$. To address this, we propose training only the tangential component of the score function when the noise scale $\sigma_{t}$ is sufficiently small, thereby avoiding the singularity associated with $\ell_{\text{quad}}^{\perp}$. To compute $\ell_{\text{quad}}^{\parallel}$, we introduce the following Tango loss,
@@
The optimal score network $s^{\parallel}_{\theta^{*}}(x,t)$ that minimizes (20) almost surely satisfies $s^{\parallel}_{\theta^{*}}(x,t)=P(x)\nabla_{x}\log p_{\sigma_{t}}(x)$. This result ensures the validity of the loss function in (20), with the proof provided in Appendix C.1.
When the noise scale is smaller than a pre-selected $c_{\mathrm{tango}}$ (i.e. $\sigma_{t}<c_{\mathrm{tango}}$), the tangential component of the score function $s^{\parallel}_{\theta}(x,t)$ is trained via the Tango loss $\ell_{\text{tango}}$. On the other hand, when $\sigma_{t}\geq c_{\mathrm{tango}}$, the singularity issues are less severe, allowing us to use the original denoising score matching loss (5) to train the entire score function $s_{\theta}(x,t)$, which includes information along the normal direction. This enables the score function to push points toward the vicinity of the manifold. The overall loss calculation is summarized in Algorithm 1.
@@ 5 Generation
Figure 1: The average error of the tangential component of the learned score function in the $x-y$ plane, along the $z$-axis and $t$-axis. From top to bottom, the plots correspond to the vanilla algorithm (Iso-DM), our proposed Niso-DM and Tango-DM, all using the rescaling technique.
We propose two sampling methods to generate samples using the learned score function $s_{\theta}$. The first method, named Reverse SDE, is based on the standard generation process in $\mathbb{R}^{n}$, with an additional projection step to ensure the samples lie on manifolds. The second method, named Annealing SDE, adopts a two-stage approach: the first stage simulates the reverse SDE, and the second stage performs annealed Langevin dynamics constrained to manifolds. The details of these methods are provided in the sections below and are summarized in Algorithms 2 and 3.
@@ 5.1 Reverse SDE
The Reverse SDE method follows the standard generation process, where samples are generated by simulating the reverse SDE (21), with $t$ going from $T$ to 00.
@@
However, the ending samples lie in the vicinity of the manifold rather than exactly on it. To address this, we add a final projection step $\pi$ to ensure that the generated points lie on the manifold.
In general, this method is not applicable to Tango-DM that involves Tango loss. Since the normal component of the score function is not learned when $t$ is small, the score function $s_{\theta}(x,t)$ does not capture information in the normal direction. For Niso-DM method, the final distribution of the reverse SDE can be expressed as $x+c_{\mathrm{niso}}N(x)\epsilon_{2}$, where $x\sim p_{0}$ and $c_{\mathrm{niso}}N(x)\epsilon_{2}$ represents noise along the normal direction. By projection onto the manifold as the final step, the samples we obtain closely approximate the target distribution.
@@ 5.2 Annealing SDE on manifolds
The Annealing SDE method adopts a two-stage sampling procedure, divided by a predefined threshold $\tilde{\sigma}$. In the first stage ($\sigma_{t}\geq\tilde{\sigma}$), we simulate the reverse SDE (21), which provides a coarse approximation of the target distribution near the manifold. In the second stage ($\sigma_{t}<\tilde{\sigma}$), we refine the samples by simulating annealed Langevin dynamics constrained to the manifold. Specifically, for a fixed $t$, we perform $n_{0}$ steps of the following Langevin dynamics:
@@
After completing $n_{0}$ steps, we decrease $t$ according to $t\leftarrow t-\Delta t$ and restart the simulation of SDE (22) using the final samples from the previous iteration as the initial states. This process is repeated until the desired level of refinement is achieved. i.e., $t=0$. This stage can be viewed as the manifold version of the Corrector algorithms in [31].
To analysis the proposed sampling method, we assume that the neural network $s_{\theta}$ has learned the optimal solution, i.e. $P(x)s_{\theta}(x,t)=\nabla_{x}^{\mathcal{M}}\log p_{\sigma_{t}}(x)$ for $x\in\mathcal{M}$. Under this assumption, the invariant distribution of the SDE (22) is $p_{\sigma_{t}}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$. According to (12) in Theorem 3.1 and (18) in Theorem 4.1, regardless of whether non-isotropic noise is used, $\nabla_{x}^{\mathcal{M}}\log p_{\sigma_{t}}(x)$ converges to $\nabla_{x}^{\mathcal{M}}\log p_{0}(x)$ as $t\to 0$. This demonstrates that by employing annealed Langevin dynamics constrained to the manifold, we can effectively sample from the target distribution $p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ on the manifold.
@@ 6 Experiments
Reversal Annealing
Iso 2.81e-4$\pm$5.66e-5 7.97e-4$\pm$6.89e-5
Niso 1.52e-4$\pm$2.60e-5 5.32e-4$\pm$3.03e-4
Tango 1.65e-4$\pm$3.91e-5 5.82e-4$\pm$1.17e-4
Iso+res 2.79e-4$\pm$9.41e-5 7.81e-4$\pm$2.69e-4
Niso+res 1.45e-4$\pm$2.68e-5 1.88e-4$\pm$4.49e-5
Tango+res 1.19e-4$\pm$1.96e-5 2.22e-4$\pm$3.03e-5
In this section, we evaluate our methods by addressing several representative problems on typical manifolds. Specifically, we consider four examples: (1) a toy model of the hyperplane in 3D space, (2) mesh manifolds in 3D space, (3) the special orthogonal group of order $10$, and (4) the level set of collective variables in the alanine dipeptide system.
Furthermore, we also consider using neural networks to approximate the normalized score functions, i.e. $s_{\theta}(x,t)=\hat{s}_{\theta}(x,t)/w_{t}$, where $\hat{s}_{\theta}(x,t)$ denotes a neural network and $w_{t}$ is the scaling factor of the optimal score function. This approach, which we refer to as the rescaling technique, is designed to improve numerical stability, as neural networks can more effectively approximate an $O(1)$ term compared to one that grows explosively. Notably, this technique is equivalent to the $\epsilon$-parameterization introduced in [26, 35].
We perform experiments with the vanilla diffusion models, as well as our proposed Niso-DM and Tango-DM, denoted as Iso, Niso, and Tango, respectively. The notation +res indicates the rescaling technique discussed above, which involves using neural networks to model the normalized score functions. After training, new samples are generated via two methods: Reverse SDE and Annealing SDE on manifolds. To evaluate the quality of the generated samples, we use several metrics to compare the generated distribution with the target distribution. We conduct our experiments using the same hyperparameters, except for those specific to each algorithm. Overall, we recommend the Niso+res training algorithm combined with the Reverse SDE generation method, as it performs best in most cases. Further details are provided in Appendix D. We also conduct ablation studies in Appendix E to analyze the bias-variance tradeoff associated with the choice of hyperparameters.
@@ 6.1 The hyperplane in 3D space
We first consider a toy model: the hyperplane embedded in a three-dimensional Euclidean space, specifically the set =$\mathcal{M}=\{(x,y,z)\in\mathbb{R}^{3}|z=0\}$. The target distribution is a mixture of Gaussian distributions with nine modes located on the plane. In this example, the score function has a closed-form solution that can be explicitly derived.
In Figure 1, we present the error of the tangential component of the score function. The results demonstrate that our proposed Niso-DM and Tango-DM significantly reduce the tangential component error, as $t$ approaches 0. Table 1 in the Appendix reports the maximum mean discrepancy (MMD), where both Niso-DM and Tango-DM show significant performance improvements.
@@ 6.2 Mesh data
We consider the Standard Bunny [32] and Spot the Cow [5], which contain meshes derived from 3D scanning a ceramic figurine of a rabbit and a cow, respectively. To create the target densities, we follow the approach in [25, 14, 4], which utilizes the eigenpairs of the Laplace-Beltrami operator on meshes. The target distribution is chosen as an equal-proportion mixture of density functions corresponding to the 0-th, 500-th, and 1000-th eigenpairs.
Table 2 reports the Jensen–Shannon (JS) divergence of the histograms on mesh faces between the generated samples and the original datasets. With the Rescaling technique, both Niso-DM and Tango-DM demonstrate significant improvements; without it, the Tango-DM still achieves improvements under the Annealing algorithm.
Table 2: Results for Stanford Bunny and Spot the Cow: JS distance of the samples generated by the Reverse SDE and Annealing SDE algorithms under different training methods. The symbol "-" indicates that the Reverse SDE sampling method is not applicable to Tango-DM. The results are reported as the mean and standard deviation over five independent runs. The notation +res indicates the rescaling technique. The bold values represent the minimum value in each column (sampling algorithm).
@@ 6.3 High-dimensional special orthogonal group
Our results demonstrate the effectiveness of the proposed method. In Table 4, we compare the sliced 1-Wasserstein [1] distances between the generated and the target distributions. These findings highlight the precision and reliability of our approach in high co-dimensional manifold settings.
@@ 6.4 Alanine dipeptide
We apply our method to alanine dipeptide, a model system frequently examined in computational chemistry. The system’s configuration is characterized by two dihedral angles, $\phi$ and $\psi$ (refer to Figure 2). The manifold consists of the configurations of the system’s 10 non-hydrogen atoms (in $\mathbb{R}^{30}$) with the angle $\phi$ fixed at $-70^{\circ}$, which represents a level set of the reaction coordinate $\phi$.
Table 4 exhibits the results of 1-Wasserstein distance between the test datasets and the generated samples, showing that our methods outperforms the vanilla models under both the Reverse SDE and the Annealing SDE sampling algorithms. However, in this case, training the normalized score functions (+res) does not lead to better results. One possible explanation is that it simultaneously rescales both the tangential and normal components, resulting in the loss of critical tangential information.
@@
Iso 1.76e-2$\pm$1.15e-2 1.86e-2$\pm$9.65e-3
Niso 5.58e-3$\pm$1.84e-3 1.16e-2$\pm$2.51e-3
Tango - 1.89e-2$\pm$2.05e-3
Iso+res 9.49e-3$\pm$1.91e-3 1.17e-2$\pm$7.29e-4
Niso+res 4.60e-3$\pm$8.24e-4 6.00e-3$\pm$7.53e-4
Tango+res - 6.42e-3$\pm$1.97e-3
Table 4: Results for dipeptide: 1-Wasserstein distance under different training methods. For more detailed explanations, refer to the caption of Table 2.

### v3

@@ 4.1 Niso-DM: Perturb data with non-isotropic noise
[…]
@@
[…]
$p_{\sigma}(\tilde{x})=(2\pi)^{-\frac{n}{2}}\int_{\mathcal{M}}p_{0}(x)(\det\Sigma_{\sigma}(x))^{-\frac{1}{2}}e^{-\frac{1}{2}(\tilde{x}-x)^{T}\Sigma_{\sigma}(x)^{-1}(\tilde{x}-x)}\mathrm{d}\sigma_{\mathcal{M}}(x),$ (16)
[…]
@@
    $\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\frac{x^{*}-\tilde{x}}{\sigma^{2\alpha}}\cdot\frac{1}{1+\sigma^{2-2\alpha}}+O(\sigma^{(1-2\alpha)\wedge 0}),$ (17)
    where the symbol $\wedge$ is defined as $a\wedge b=\min\{a,b\}$.
@@
Eq. (17) shows that, by introducing additional noise along the normal direction, the scale of the normal component is reduced from $O(1/\sigma^{2})$ to $O(1/\sigma^{2\alpha})$. Eq. (18) states that this approach does not affect the characterization of the distribution on the manifold (i.e., the tangential component). As the noise level parameter $\sigma_{t}$ approaches zero, the score function of the perturbed density $p_{\sigma_{t}}$ on the manifold converges to the Riemannian score function of $p_{0}$. Therefore, by employing non-isotropic noise, we can still generate samples that adhere to the original distribution of the dataset.
@@ 4.2 Tango-DM: Learn only the tangential component of the score function
Recall that the projection matrix $P(x)$ in (8) is well-defined in the ambient space $\mathbb{R}^{n}$. Accordingly, we define $s^{\parallel}_{\theta}(x,t):=P(x)s_{\theta}(x,t)$ for $x\in\mathbb{R}^{n}$, which represents the tangential component of the parametrized vector field.
As discussed in Section 3 (see (13)), the loss function (4) and (5) can be decomposed into two parts, $\ell_{\text{quad}}^{\parallel}$ and $\ell_{\text{quad}}^{\perp}$, and the singularity issue is associated with the part $\ell_{\text{quad}}^{\perp}$. To address this, we propose training only the tangential component of the score function using the loss $\ell_{\text{quad}}^{\parallel}$ when the noise scale $\sigma_{t}$ is sufficiently small, thereby avoiding the singularity associated with $\ell_{\text{quad}}^{\perp}$. To compute $\ell_{\text{quad}}^{\parallel}$, we introduce the following Tango loss
@@
The optimal score network $s^{\parallel}_{\theta^{*}}(x,t)$ that minimizes (20) satisfies $s^{\parallel}_{\theta^{*}}(x,t)=P(x)\nabla_{x}\log p_{\sigma_{t}}(x)$ (See Appendix C.1 for the proof). This result ensures the validity of the loss function in (20).
When the noise scale is smaller than a pre-selected $c_{\mathrm{tango}}$ (i.e. $\sigma_{t}<c_{\mathrm{tango}}$), the tangential component of the score function $s^{\parallel}_{\theta}(x,t)$ is trained via the Tango loss $\ell_{\text{tango}}$. On the other hand, when $\sigma_{t}\geq c_{\mathrm{tango}}$, the singularity issue is less severe, allowing us to use the original denoising score matching loss (5) to train the entire score function $s_{\theta}(x,t)$, which includes information along the normal direction. This enables the score function to push points toward the vicinity of the manifold. The overall loss calculation is summarized in Algorithm 1 in the Appendix.
@@ 5 Generation
Figure 1: The average error of the tangential component of the learned score function in the xyx-y plane, along the $z$-axis and $t$-axis, with $\sigma_{\min}=0.01$. From top to bottom, the plots correspond to the vanilla algorithm (Iso-DM), our proposed Niso-DM and Tango-DM, all using the rescaling technique.
We propose two sampling methods to generate samples using the learned score function $s_{\theta}$. The first method, named Reverse SDE, is based on the standard generation process in $\mathbb{R}^{n}$, with an additional projection step to ensure that the samples lie on manifolds. The second method, named Annealing SDE, adopts a two-stage approach, where the first stage simulates the standard reverse SDE, and the second stage performs annealed Langevin dynamics constrained to manifolds. The details of these methods are provided in the sections below and are summarized in Algorithms 2 and 3 in the Appendix.
@@ 5.1 Reverse SDE
The first method follows the standard generation process, where samples are generated by simulating the reverse SDE (21) with $t$ decreasing from $T$ to 0, which is derived from (2) by setting the drift term $f$ to 0.
@@
However, the ending samples lie in the vicinity of the manifold rather than exactly on it. To address this, we add a final projection step to ensure that the generated points lie on the manifold.
In general, this method is not applicable to Tango-DM, which utilizes the Tango loss in (20). Since the normal component of the score function is not learned when $t$ is small, the score function $s_{\theta}(x,t)$ does not capture information in the normal direction. For Niso-DM, the final distribution of the reverse SDE is the same as the distribution of $x+c_{\mathrm{niso}}N(x)\epsilon_{2}$, where $x\sim p_{0}$ and $c_{\mathrm{niso}}N(x)\epsilon_{2}$ represents noise along the normal direction. By employing a projection onto the manifold in the subsequent step, the obtained samples can accurately approximate the target distribution.
@@ 5.2 Annealing SDE on manifolds
The second method adopts a two-stage sampling procedure with a predefined threshold $\tilde{\sigma}$. In the first stage ($\sigma_{t}\geq\tilde{\sigma}$), we simulate the reverse SDE (21), which provides a coarse approximation of the target distribution. In the second stage ($\sigma_{t}<\tilde{\sigma}$), we refine the samples by simulating annealed Langevin dynamics constrained to the manifold. Specifically, for a fixed $t$, we perform $n_{0}$ steps of the following Langevin dynamics:
@@
After that, we decrease $t$ according to $t\leftarrow t-\Delta t$ and restart the simulation of SDE (22) using the final samples from the previous iteration as the initial states. This process is repeated until the desired level of refinement is achieved. i.e., $t=0$. This stage can be viewed as the manifold version of the Corrector algorithms in [38].
To analysis the proposed sampling method, we assume that the neural network $s_{\theta}$ has learned the optimal solution, i.e. $P(x)s_{\theta}(x,t)=\nabla_{x}^{\mathcal{M}}\log p_{\sigma_{t}}(x)$ for $x\in\mathcal{M}$. Under this assumption, the invariant distribution of the SDE (22) is $p_{\sigma_{t}}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$. According to (12) in Theorem 3.1 and (18) in Theorem 4.1, regardless of whether non-isotropic noise is used, $\nabla_{x}^{\mathcal{M}}\log p_{\sigma_{t}}(x)$ converges to $\nabla_{x}^{\mathcal{M}}\log p_{0}(x)$ as $t\to 0$. This demonstrates that, by employing annealed Langevin dynamics constrained to the manifold, we can effectively sample from the target distribution $p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$ on the manifold.
@@ 6 Experiments
In this section, we evaluate our methods by studying several representative problems on typical manifolds. Specifically, we consider four datasets on four different types of submanifolds: (1) a hyperplane in 3D space, (2) mesh manifolds in 3D space, (3) the special orthogonal group of order $10$, and (4) the level set of dihedral angles in the molecular system alanine dipeptide. Our code is available at: https://github.com/ZichenLiu1999/NisoTangoDM.
We perform experiments with the vanilla diffusion models, as well as our proposed Niso-DM and Tango-DM, denoted as Iso, Niso, and Tango, respectively. Moreover, we consider using neural networks to approximate normalized score functions, defined as $s_{\theta}(x,t)=\hat{s}_{\theta}(x,t)/w_{t}$, where $\hat{s}_{\theta}(x,t)$ is a neural network and $w_{t}$ is the scaling factor of the optimal score function. This rescaling technique improves numerical stability by allowing neural networks to approximate an $O(1)$ term instead of one that grows explosively. Notably, this technique is equivalent to the $\epsilon$-parameterization introduced in [33, 47]. We denote this method with the suffix +res. After training, new samples are generated via two methods: Reverse SDE and Annealing SDE on manifolds. Although we use "vanilla" to refer to the Iso and Iso+res methods, these two methods also involve manifold information as we utilize the projection operator during the generation process.
To evaluate the quality of the generated samples, we use several metrics to compare the generated distribution with the target distribution. The choice of metrics balances the ability to capture on-manifold distribution differences with reasonable computational cost. We conduct our experiments using the same hyperparameters, except for those specific to each algorithm. Overall, we recommend the Niso+res training algorithm combined with the Reverse SDE generation method, as it performs best in most cases. In experiments on mesh data and the special orthogonal group, we also conduct Riemannian sliced score matching (RSSM, for simplicity) in [8, 17] as a baseline. Further experimental details are provided in Appendix D. We also conduct ablation studies in Appendix E to analyze the bias-variance tradeoff associated with the choice of hyperparameters.
@@ 6.1 The hyperplane in 3D space
We first consider a toy model: the hyperplane embedded in a three-dimensional Euclidean space, specifically the set $\mathcal{M}=\{(x,y,z)\in\mathbb{R}^{3}|z=0\}$. The target distribution is a mixture of Gaussian distributions with nine modes located on the plane. In this example, the score function has a closed-form solution that can be explicitly derived.
In Figure 1, we present the error of the tangential component of the learned score function. The results demonstrate that our proposed Niso-DM and Tango-DM significantly reduce the error of the tangential component, as $t$ approaches 0. Table 5 in the Appendix reports the maximum mean discrepancy (MMD), where both Niso-DM and Tango-DM show significant performance improvements.
@@ 6.2 Mesh data
We consider the Stanford Bunny [40] and Spot the Cow [7], which contain meshes derived from 3D scanning ceramic figurines of a rabbit and a cow, respectively. To create the target densities, we follow the approach in [32, 19, 6], which utilizes the eigenpairs of the Laplace-Beltrami operator on meshes. The target distribution is chosen as an equal-proportion mixture of density functions corresponding to the 0-th, 500-th, and 1000-th eigenpairs. The manifold $\mathcal{M}$ is a polyhedron composed of triangular faces, differing from the definition provided in Section 2.2. Although the manifold is not smooth along the shared edges between adjacent faces, which form a zero-measure set, numerical performance ensures that this is an acceptable setup.
Table 1 reports the Jensen–Shannon (JS) divergence of the histograms on mesh faces between the generated samples and the original datasets. With the Rescaling technique, both Niso-DM and Tango-DM demonstrate significant improvements; without it, Tango-DM still achieves improvements under the Annealing algorithm. Furthermore, we suspect that the suboptimal performance of RSSM arises from the challenges in achieving a uniform distribution on manifolds with complex geometric structures, such as the narrow neck in "Spot the Cow."
Table 1: Results for the mesh data: JS divergence of the samples generated by the Reverse SDE and Annealing SDE algorithms under different training methods. The symbol "-" indicates that the Reverse SDE sampling method is not applicable to Tango-DM. The results are reported as the mean and standard deviation over five independent runs. The suffix +res indicates the rescaling technique.
@@ 6.3 High-dimensional special orthogonal group
Our results demonstrate the effectiveness of the proposed methods. In Table 3, we compare the sliced 1-Wasserstein [2] distances between the generated and the target distributions. These results highlight the accuracy and reliability of our approach in high dimensional manifold settings. Moreover, our approach outperforms RSSM, while the vanilla algorithm does not.
@@ 6.4 Alanine dipeptide
We apply our method to alanine dipeptide, a model system frequently examined in computational chemistry. The system’s configuration is characterized by two dihedral angles, $\phi$ and $\psi$ (refer to Figure 4). The manifold consists of the configurations of the system’s 10 non-hydrogen atoms (in $\mathbb{R}^{30}$) with the angle $\phi$ fixed at $70^{\circ}$, which is a level set of the dihedral angle $\phi$.
Table 3 exhibits the results of the 2-Wasserstein distance between the test datasets and the generated samples, showing that our methods outperform the vanilla models under both the Reverse SDE and the Annealing SDE sampling algorithms. However, in this case, training the normalized score functions (+res) does not lead to better results. One possible explanation is that it simultaneously rescales both the tangential and normal components, resulting in the loss of critical tangential information.
@@
[…]
Tango - 1.89e-$\pm2$.05e-3
[…]
Table 3: Results for dipeptide: 2-Wasserstein distance under different training methods. For more detailed explanations, refer to the caption of Table 1.